Newton's Method with two variables

This problem focuses on the system of simultaneous equations

x^{x}+y^{y}=1000

x^{y}+y^{x}=100

Choose any initial point (x_{0},y_{0}) anywhere near either solution (x_{*},y_{*}) of the system of equations

For this problem you need to only complete at least one iteration starting with numerical values for (x_{0},y_{0}) terminating with numerical values for (x_{1},y_{1})
Answer #1

→ Given, **+ yt = 1000 x² + y² = 100 Newtons iteration method f ,y) =o 718,4)=0 Let Crosto) be initial approximation. Then hax oy fx = Of(0,1) = x*Cl+logu) fyer Cany) = y*(1+ logy) In = dg (x,y) = yoxt, ye logy og (vy)= .y*4 2 y 10g * Let Crouto) bDx = -972 -96 = -6017.88 6-29 Dy = | 56.66 -972 | .-4467136 | - | na noge -(1985) = 16.95 Ks - - -/-4457.5) - 12:58 29 = Xot